A 7.15 cm-diameter circular loop of wire is in a 1.11 T magnetic field. The loop is removed from the field in 0.190 s. What is the average induced emf?
(in V)

A: 1.71
Answer #1

induced emf is defiend as the rate of change of magnetic flux

in mathematical form,

induced emf e = NAdB/dt

where A = area

N = no. of tunrs

dB/dt = rate of change of amgnetic flux

so here emf e = 1 * 3.14* (0.0715/2)^2 * 1.11/0.19

emf e = 2.35
Answer #2

emf = |dflux/dt|
= |d(BA)/dt|
= A * |dB/dt|
= pi*0.0357^2 * | (0 - 1.10) / 0.19 |
= 2.3 * 10^ -2 V

option C
Answer #3

Area of loop=pi*d2/4=3.14*(0.0715)2/4=4.015*10-3 m2

Magnetic field B=1.11 T

time t=0.190 s

Emf=Field*Area/time

                =4.015*10-3*1.11/0.190=0.0235 V

Or, C. 2.35*10-2 V
Answer #4

emf generated = -\frac{\mathrm{d} \phi }{\mathrm{d} t}=-\frac{BA-0}{t}=2.35*10^{-2}V